OpenStudy (anonymous):

Convert each number into an exponential number. 100,000 = 1000 = 343 = 16 = 81 = 3125 =

OpenStudy (rogue):

10^n = 1 with n zeros, so 100,000 = 10^5. 1,000 = 10^3 343 = 7^3 16 = 4^2 = 2^4 81 = 9^2 = 3^4 3125 = 5^5

OpenStudy (rogue):

The more you practice with squares, cubes, and roots, the easier this becomes =)